# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v1

@@ 4.1. Grigor’yan Parabolic Maximum Principle
We borrow the following version of Grigor’yan parabolic maximum principle from [Gri09, Theorem 11.9]. This result is originally designed to obtain the uniqueness of the bounded solution to the Cauchy problem, which helps to prove the stochastic completeness of weighted manifolds. We refer the proof of Theorem 4.1 and some related discussions to [Gri06, Gri09].
@@
Here $f$ is some bounded function and the diffusion operator is given by $\Delta_{f}\triangleq\Delta+\langle\nabla f,\nabla\rangle$. Assume that for some $x_{0}\in M$ and for all $r>0$,
@@
By examining the proof of Theorem 4.1, we notice, as also shown in Y. Li [Li15], that the result has the following variant, which we will mainly refer to later.
@@
For completeness, we explain here why Theorem 4.2 is valid. Indeed, since $\phi(x)=x_{+}$ is non-decreasing, convex, continuous and piece-wise smooth, the condition (4.6) yields that
@@
Also in Theorem 4.1 the result still holds for $F$ satisfying $\partial_{t}F\leq\Delta_{f}F$ and $F\geq 0$, since the only step using the PDE itself is the equation [Gri09, (11.37)], which turns into an inequality in the new case and still yields the following steps. Hence we apply Theorem 4.1 to $F_{+}$ and obtain Theorem 4.2. Furthermore, one is easy to check from the original proof that replacing $\Delta$ by $\sigma\Delta$ in the definition of $\Delta_{f}$ will not change the result as long as $\sigma>0$. In our case, it is convenient to choose that
@@
is bounded and satisfies that $K=\nabla h$.
@@ 4.2. Logarithmic Gradient Estimate
Now we turn to the proof of our logarithmic estimates Proposition 4.1. We shall first focus on the gradient estimate (4.1) in this subsection. First we provide a technical lemma concerning the propagation results of some quantities.
@@
$\displaystyle(\partial_{t}-\Delta_{f})\Big{(}\frac{|\nabla g|^{2}}{g}\Big{)}=-\frac{2\sigma}{g}\Big{|}\partial_{ij}g-\frac{\partial_{i}g\partial_{j}g}{g}\Big{|}^{2}-\frac{2}{g}\nabla g\cdot\nabla(K\ast\omega)\cdot\nabla g\leq\frac{C}{1+t}\frac{|\nabla g|^{2}}{g},$ (4.9)
$\displaystyle(\partial_{t}-\Delta_{f})(g\log g)=-\sigma\frac{|\nabla g|^{2}}{g},$ (4.10)
The proof of Lemma 4.1 is nothing but technical calculations, hence we omit here and refer to [FW23, Section 6] for details.
Now for the short-time result, we construct the first auxiliary function to be propagated as
$F(t,x)=\frac{|\nabla g|^{2}}{g}+\frac{C}{\sigma}g\log g-C_{1}g.$
From Lemma 4.1 above we have the propagation of the auxiliary function reads
$(\partial_{t}-\Delta_{f})F\leq C\frac{|\nabla g|^{2}}{g}-C\frac{|\nabla g|^{2}}{g}.$
The initial value of $F$ simply reads
$F(0,x)=\frac{|\nabla g_{0}|^{2}}{g_{0}}+\frac{C}{\sigma}g_{0}\log g_{0}-C_{1}g_{0}.$
Hence $F(0,\cdot)\leq 0$ is equivalent to
$|\nabla\log g_{0}|^{2}+\frac{C}{\sigma}\log g_{0}\leq C_{1},$
which is valid thanks to the initial assumptions of $\nabla\log g_{0}$ and the Gaussian decay from above of $g_{0}$. Now we are only left to check the growth assumption (4.7). Recall that $\nabla g\in L^{\infty}$ and $g\in L^{\infty}$ from Lemma 3.4, and from the Gaussian lower bound Lemma 3.3
$g(x,t)\geq\frac{1}{C(1+t)^{C}}\exp\Big{(}-C\frac{|x|^{2}}{1+t}\Big{)}.$
Consequently we check the growth assumption by
@@
$|\nabla\log g|^{2}+C\log g\leq C_{1}.$ (4.11)
Recall from the Gaussian lower bound Lemma 3.3 that
$\log g\geq-C-C\log(1+t)-C\frac{|x|^{2}}{1+t}.$
Substituting that into (4.11) gives the gradient estimate (4.1) for short time.
In order to obtain the gradient estimate bound (4.1) for a long time, we need to construct another auxiliary function, say
$F(t,x)=\phi\frac{|\nabla g|^{2}}{g}+g\log g-C_{1}g,$
for some $\phi(t)$ to be determined. We deduce that the propagation of the new auxiliary function reads
$(\partial_{t}-\Delta_{f})F\leq\Big{(}\phi^{\prime}+\frac{C}{t}\phi-\sigma\Big{)}\frac{|\nabla g|^{2}}{g}.$
Hence if we choose $C_{1}$ large and $\phi$ satisfying
$\phi^{\prime}+\frac{C}{t}\phi-\sigma=0,\quad\phi(0)=0,$
then the initial assumption (4.6) holds immediately. Obviously $\phi(t)=\sigma t/(C+1)$ works. The rest of assumptions of Theorem 4.2 follow as in the short time case. Hence we arrive at $F\leq 0$, or
$|\nabla\log g|^{2}+\frac{C+1}{\sigma t}\log g\leq\frac{C_{1}(C+1)}{\sigma t}.$ (4.12)
Recall from the Gaussian lower bound Lemma 3.3 that
$\log g\geq-C-C\log(1+t)-C\frac{|x|^{2}}{1+t}.$
Substituting that into (4.12) gives the gradient estimate (4.1) for long time.
@@ 4.3. Logarithmic Hessian Estimate
$\displaystyle(\partial_{t}-\Delta_{f})\Big{(}\frac{|\nabla^{2}g|^{2}}{g}\Big{)}\leq\frac{C}{1+t}\frac{|\nabla^{2}g|^{2}}{g}+\frac{C}{(1+t)^{2}}\frac{|\nabla g|^{2}}{g}.$ (4.13)
$\displaystyle(\partial_{t}-\Delta_{f})\Big{(}g(\log g)^{2}\Big{)}=-\frac{2\sigma}{g}|\nabla g|^{2}(1+\log g).$ (4.14)
The proof of Lemma 4.2 is also nothing but technical calculations, hence we also omit here and refer to [FW23, Section 7] for details.
Now for the short-time result, we construct the first auxiliary function to be propagated as
$F(t,x)=e^{-Ct}\frac{|\nabla^{2}g|^{2}}{g}-Cg(\log g)^{2}+C_{1}g\log g-C_{2}g$
where $C_{1}\geq 2C\sigma(1+L)$ for some $L>0$ satisfying $\log g\leq L$ and $C_{2}$ large. We deduce from Lemma 4.2 that the propagation of the auxiliary function reads
$(\partial_{t}-\Delta_{f})F\leq Ce^{-Ct}\frac{|\nabla^{2}g|^{2}}{g}+Ce^{-Ct}\frac{|\nabla g|^{2}}{g}-Ce^{-Ct}\frac{|\nabla^{2}g|^{2}}{g}+2C\sigma\frac{|\nabla g|^{2}}{g}(1+\log g)-C_{1}\sigma\frac{|\nabla g|^{2}}{g}\leq 0.$
Also, the initial assumption (4.6) $F(0,\cdot)\leq 0$ is equivalent to
$\frac{|\nabla^{2}g_{0}|^{2}}{g_{0}^{2}}+2C_{1}\log g_{0}\leq C_{2}+C(\log g_{0})^{2}.$
By the initial assumption on $\nabla^{2}\log g_{0}$, we have the left-handside bounded by some $C_{3}+C_{3}|x|^{4}$ from above; while by the Gaussian upper and lower bound Lemma 3.2 and Lemma 3.3, we have the right-handside bounded by some $C_{4}+C_{4}|x|^{4}$ from below for $|x|$ large, and hence for all $x$. Here $C_{4}$ can be arbitrarily large by choosing $C,C_{2}$ large. Hence the initial condition is satisfied.
Now we are only left to check the growth assumption (4.7). Recall that $\nabla^{2}g\in L^{\infty}$, $\nabla g\in L^{\infty}$ and $g\in L^{\infty}$ from Lemma 3.4, and from the Gaussian lower bound Lemma 3.3 that
$g(x,t)\geq\frac{1}{C(1+t)^{C}}\exp\Big{(}-C\frac{|x|^{2}}{1+t}\Big{)}.$
Those imply that
@@
$\frac{|\nabla^{2}g|^{2}}{g^{2}}\leq e^{Ct}\Big{(}C(\log g)^{2}-C\log g+C\Big{)}.$ (4.15)
Recall from the Gaussian upper and lower bound Lemma 3.2 and Lemma 3.3 that
$\log g\geq-C-C\log(1+t)-C\frac{|x|^{2}}{1+t},\quad\log g\leq C-\log(1+t)-\frac{|x|^{2}}{8t+C}.$
Substituting those into (4.15) gives the Hessian estimate (4.2) for short time.
@@
In order to obtain the Hessian estimate bound (4.2) for a long time, we need to make use of the non-positive term which is abandoned in (4.9) when discussing the logarithmic gradient estimate, and construct another auxiliary function. Rewrite (4.9) as
$(\partial_{t}-\Delta_{f})\Big{(}\frac{|\nabla g|^{2}}{g}\Big{)}\leq-\frac{|\nabla^{2}g|^{2}}{g}+2\frac{|\nabla g|^{4}}{g^{3}}+\frac{C}{1+t}\frac{|\nabla g|^{2}}{g},$ (4.16)
and define our new auxiliary function as
$F(x,t)=\phi\frac{|\nabla^{2}g|^{2}}{g}+\psi\frac{|\nabla g|^{2}}{g}-A_{1}g(\log g)^{2}+A_{2}g\log g-A_{3}g,$
@@
for some $\phi(t)$ and $\psi(t)$ to be determined. We deduce that the propagation of our new auxiliary function reads
$\displaystyle(\partial_{t}-\Delta_{f})F\leq$ $\displaystyle\Big{(}\phi^{\prime}+\frac{C}{t}\phi-\psi\Big{)}\frac{|\nabla^{2}g|^{2}}{g}$
$\displaystyle+\Big{(}\frac{C}{t^{2}}\phi+\psi^{\prime}+\frac{C}{t}\psi+2\psi|\nabla\log g|^{2}+2\sigma A_{1}\log g+2\sigma A_{1}-\sigma A_{2}\Big{)}\frac{|\nabla g|^{2}}{g}.$
Also we set
$\phi(0)=\psi(0)=0.$
Recall the proof of the gradient estimate, we obtain for $\psi(t)=A_{1}t/(C+1)$ that
$2\psi|\nabla\log g|^{2}+2A_{1}\log g\leq C_{1}A_{1}(C+1).$
Now pick $\phi(t)=t^{2}/\Big{(}(C+1)(C+2)\Big{)}$, we are able to cancel the first term on the right-hand side. Pick $A_{2}$ large to cancel the second term. The initial condition (4.6) is also satisfied for large $A_{3}$. The growth assumption (4.7) is verified similarly as the short time case. Hence we arrive at $F\leq 0$, or
$t^{2}\frac{|\nabla^{2}g|^{2}}{g^{2}}\leq C(\log g)^{2}+C.$ (4.17)
Recall from the Gaussian upper and lower bound Lemma 3.2 and Lemma 3.3 that
$\log g\geq-C-C\log(1+t)-C\frac{|x|^{2}}{1+t}\text{ and }\log g\leq C-\log(1+t)-\frac{|x|^{2}}{8t+C}.$
Substituting those into (4.17) gives the Hessian estimate (4.2) for long time.
@@
Acknowledgments. X. F. and Z. W. are partially supported by the National Key R&D Program of China, Project Number 2021YFA1002800, NSFC grant No.12171009, Young Elite Scientist Sponsorship Program by China Association for Science and Technology (CAST) No. YESS20200028 and the Fundamental Research Funds for the Central Universities (the start-up fund), Peking University.

### v2

@@ 4.1. Grigor’yan Parabolic Maximum Principle
We borrow the following version of Grigor’yan parabolic maximum principle from [20, Theorem 11.9]. This result is originally designed to obtain the uniqueness of the bounded solution to the Cauchy problem, which helps to prove the stochastic completeness of weighted manifolds. We refer to [21, 20] for the proof of Theorem 4.1 and some related discussions.
@@
Here $f$ is some bounded function, and the diffusion operator is given by $\Delta_{f}\triangleq\Delta+\langle\nabla f,\nabla\rangle$. Assume that for some $x_{0}\in M$ and for all $r>0$,
@@
By examining the proof of Theorem 4.1, we notice, as also shown in Y. Li [37], that the result has the following variant, which we will mainly refer to later.
@@
For the sake of completeness, we explain here why Theorem 4.2 is valid. Indeed, since $\phi(x)=x_{+}$ is non-decreasing, convex, continuous and piece-wise smooth, the condition (4.8) yields that
@@
Also in Theorem 4.1 the result still holds for $F$ satisfying $\partial_{t}F\leq\Delta_{f}F$ and $F\geq 0$, since the only step using the PDE itself is the equation [20, (11.37)], which turns into an inequality in the new case and still yields the following steps. Therefore, we apply Theorem 4.1 to $F_{+}$ and obtain Theorem 4.2. Furthermore, it is straightforward to check from the original proof that replacing $\Delta$ by $\sigma\Delta$ in the definition of $\Delta_{f}$ will not change the result as long as $\sigma>0$. In our case, it is convenient to choose
@@
is bounded and satisfies $K=\nabla h$.
@@ 4.2. Logarithmic Gradient Estimate
Now we turn to the proof of our logarithmic estimates, i.e. Proposition 4.1. In this subsection, we focus on the gradient estimate (4.1), and the Hessian estimate (4.2) will be discussed in the next subsection.
First, we provide a technical lemma concerning the propagation results of some quantities.
@@
$\displaystyle(\partial_{t}-\Delta_{f})\Big{(}\frac{|\nabla g|^{2}}{g}\Big{)}=-\frac{2\sigma}{g}\Big{|}\partial_{ij}g-\frac{\partial_{i}g\partial_{j}g}{g}\Big{|}^{2}-\frac{2}{g}\nabla g\cdot\nabla(K\ast\omega)\cdot\nabla g\leq\frac{2M_{1}}{1+t}\frac{|\nabla g|^{2}}{g},$ (4.11)
$\displaystyle(\partial_{t}-\Delta_{f})(g\log g)=-\sigma\frac{|\nabla g|^{2}}{g}.$ (4.12)
We recall that the constant $M_{1}$ is defined in Lemma 3.1.
The proof of Lemma 4.1 is nothing but technical computations, which we omit here and refer to [16, Section 6] for details.
Short Time Estimate: We construct the first auxiliary function to be propagated:
$F(t,x)=\frac{|\nabla g|^{2}}{g}+\frac{C_{1}}{\sigma}g\log g-C_{2}g$
with some constants $C_{1},C_{2}$ to be determined. From Lemma 4.1 above, the propagation of the auxiliary function reads
$(\partial_{t}-\Delta_{f})F\leq\frac{2M_{1}}{1+t}\frac{|\nabla g|^{2}}{g}-C_{1}\frac{|\nabla g|^{2}}{g}\leq 0$
as long as $C_{1}\geq 2M_{1}$. The initial value of $F$ simply reads
$F(0,x)=\frac{|\nabla g_{0}|^{2}}{g_{0}}+\frac{C_{1}}{\sigma}g_{0}\log g_{0}-C_{2}g_{0}.$
Hence, $F(0,\cdot)\leq 0$ is equivalent to
$|\nabla\log g_{0}|^{2}+\frac{C_{1}}{\sigma}\log g_{0}\leq C_{2}.$
We recall the initial assumptions:
$|\nabla\log g_{0}(m,x)|\lesssim 1+|x|,\quad g_{0}(m,x)\leq C_{0}\exp(-C_{0}^{-1}|x|^{2}),$
which allow us to choose $C_{1}$ large enough to eliminate the spatial growth in $|\nabla\log g_{0}|^{2}$ and then choose $C_{2}$ large enough to conclude that $F(0,\cdot)\leq 0$. Now we only have to check the growth assumption (4.9). Recall that $\nabla g\in L^{\infty}$ and $g\in L^{\infty}$ from Lemma 3.4, and that
$g(x,t)\geq\frac{1}{C(1+t)}\exp\Big{(}-C\frac{|x|^{2}}{1+t}\Big{)}$
from the Gaussian lower bound Lemma 3.3. Consequently, we check the growth assumption by
@@
$|\nabla\log g|^{2}+\frac{C_{1}}{\sigma}\log g\leq C_{2}.$ (4.13)
Long Time Estimate: In order to obtain the logarithmic gradient estimate (4.1) with optimal time dependence, we need a small trick motivated by [23]. For any fixed time $\tau>0$, we define the following time-rescaling function $\tilde{g}$ by pushing forward the time variable by $\tau/2$:
$\tilde{g}(t,m,x)=g(t+\frac{\tau}{2},m,x),$
which also solves the same nonlinear Fokker-Planck equation (1.7) with the initial data changing into $\tilde{g}(0,\cdot)=g(\frac{\tau}{2},\cdot)$. Hence the corresponding asymptotic decay estimates read as
$\|\nabla^{k}\tilde{g}\|_{L^{\infty}}\leq M_{k}(1+\frac{\tau}{2}+t)^{-1-\frac{k}{2}},\quad\|\nabla^{k}K\ast\tilde{g}\|_{L^{\infty}}\leq M_{k}(1+\frac{\tau}{2}+t)^{-\frac{1+k}{2}}.$
We also need to construct another auxiliary function to be propagated:
$F(t,x)=\phi\frac{|\nabla\tilde{g}|^{2}}{\tilde{g}}+\tilde{g}\log\tilde{g}-C_{4}\tilde{g},$
for some $\phi(t)$ and constant $C_{4}$ to be determined. We deduce that the propagation of the new auxiliary function reads
$(\partial_{t}-\Delta_{f})F\leq\Big{(}\phi^{\prime}+\frac{2M_{1}}{1+t}\phi-\sigma\Big{)}\frac{|\nabla\tilde{g}|^{2}}{\tilde{g}}.$
Hence, if we choose $C_{4}=\log\frac{2M_{0}}{\tau}$ and $\phi$ satisfying
$\phi^{\prime}+\frac{2M_{1}}{1+t}\phi-\sigma\leq 0,\quad\phi(0)=0,$
then the initial assumption (4.8) is satisfied immediately. Obviously $\phi(t)=\frac{\sigma t}{2M_{1}+1}$ works. The rest of the assumptions of Theorem 4.2 follow as in the short time case. Therefore, we arrive at $F\leq 0$, or
$|\nabla\log\tilde{g}|^{2}+\frac{2M_{1}+1}{\sigma t}\log\tilde{g}\leq\frac{C_{4}(2M_{1}+1)}{\sigma t}.$
Take $t=\tau/2$ and return to the original function $g$:
$|\nabla\log g(\tau,m,x)|^{2}+\frac{4M_{1}+2}{\sigma\tau}\log g(\tau,m,x)\leq\frac{(4M_{1}+2)(\log 2M_{0}-\log\tau)}{\sigma\tau}.$ (4.14)
Combining (4.13) for $t\leq 1$ and (4.14) for $t\geq 1$, we can choose some appropriate large constants $B_{1},B_{2}$ to conclude the logarithmic gradient estimate (4.1).
@@ 4.3. Logarithmic Hessian Estimate
$\displaystyle(\partial_{t}-\Delta_{f})\Big{(}\frac{|\nabla^{2}g|^{2}}{g}\Big{)}\leq\frac{4M_{1}+M_{2}}{1+t}\frac{|\nabla^{2}g|^{2}}{g}+\frac{M_{2}}{(1+t)^{2}}\frac{|\nabla g|^{2}}{g}.$ (4.15)
$\displaystyle(\partial_{t}-\Delta_{f})\Big{(}g(\log g)^{2}\Big{)}=-2\sigma\frac{|\nabla g|^{2}}{g}(1+\log g).$ (4.16)
We recall that $M_{1},M_{2}$ are defined in Lemma 3.1.
The proof of Lemma 4.2 is also nothing but technical computations, which we also omit here and refer to [16, Section 7] for details.
Short Time Estimate: We construct the first auxiliary function to be propagated:
$F(t,x)=e^{-(4M_{1}+M_{2})t}\frac{|\nabla^{2}g|^{2}}{g}-C_{5}g(\log g)^{2}+C_{6}g\log g-C_{7}g$
with some constants $C_{5},C_{6},C_{7}$ to be determined. From Lemma 4.2 above, the propagation of the auxiliary function reads
$\displaystyle(\partial_{t}-\Delta_{f})F\leq$ $\displaystyle\,-(4M_{1}+M_{2})e^{-(4M_{1}+M_{2})t}\frac{|\nabla^{2}g|^{2}}{g}$
$\displaystyle\,+\frac{4M_{1}+M_{2}}{1+t}e^{-(4M_{1}+M_{2})t}\frac{|\nabla^{2}g|^{2}}{g}+\frac{M_{2}}{(1+t)^{2}}e^{-(4M_{1}+M_{2})t}\frac{|\nabla g|^{2}}{g}$
$\displaystyle\,+2C_{5}\sigma\frac{|\nabla g|^{2}}{g}(1+\log g)-C_{6}\sigma\frac{|\nabla g|^{2}}{g}$
[…]
The initial value of $F$ simply reads
$F(0,x)=\frac{|\nabla^{2}g_{0}|^{2}}{g_{0}}-C_{5}g_{0}(\log g_{0})^{2}+C_{6}g_{0}\log g_{0}-C_{7}g_{0}.$
Hence, $F(0,\cdot)\leq 0$ is equivalent to
$\frac{|\nabla^{2}g_{0}|^{2}}{g_{0}^{2}}+C_{6}\log g_{0}\leq C_{7}+C_{5}(\log g_{0})^{2}.$
We recall the initial assumptions:
$|\nabla\log g_{0}(m,x)|^{2},|\nabla^{2}\log g_{0}(m,x)|\lesssim 1+|x|^{2},\quad g_{0}(m,x)\leq C_{0}\exp(-C_{0}^{-1}|x|^{2}),$
which allow us to choose $C_{5}$ large enough to eliminate the spatial growth in $|\nabla^{2}g_{0}|^{2}/g_{0}^{2}$, and then choose $C_{6}$ large enough to satisfy that $(\partial_{t}-\Delta_{f})F\leq 0$, and finally choose $C_{7}$ large enough to conclude that $F(0,\cdot)\leq 0$. Now we only have to check the growth assumption (4.9). Recall that $\nabla^{2}g\in L^{\infty}$, $\nabla g\in L^{\infty}$ and $g\in L^{\infty}$ from Lemma 3.4, and that
$g(x,t)\geq\frac{1}{C(1+t)}\exp\Big{(}-C\frac{|x|^{2}}{1+t}\Big{)}$
from the Gaussian lower bound Lemma 3.3. Consequently, we check the growth assumption by
@@
$\frac{|\nabla^{2}g|^{2}}{g^{2}}\leq e^{(4M_{1}+M_{2})t}\Big{(}C_{5}(\log g)^{2}-C_{6}\log g+C_{7}\Big{)}.$ (4.17)
@@
Long Time Estimate: In order to obtain the logarithmic Hessian estimate (4.2) with optimal time dependence, we also need the small trick motivated by [23]. For any fixed time $\tau>0$, we define the time-rescaling function $\tilde{g}$ as in the previous subsection. We also need to make use of the non-positive term which is abandoned in (4.11) when discussing the logarithmic gradient estimate. Rewrite (4.11) into
$(\partial_{t}-\Delta_{f})\Big{(}\frac{|\nabla g|^{2}}{g}\Big{)}\leq-\sigma\frac{|\nabla^{2}g|^{2}}{g}+2\sigma\frac{|\nabla g|^{4}}{g^{3}}+\frac{2M_{1}}{1+t}\frac{|\nabla g|^{2}}{g}.$ (4.18)
We construct our new auxiliary function to be propagated:
$F(t,x)=\phi\frac{|\nabla^{2}\tilde{g}|^{2}}{\tilde{g}}+\psi\frac{|\nabla\tilde{g}|^{2}}{\tilde{g}}-C_{8}\tilde{g}(\log\tilde{g})^{2}+C_{9}\tilde{g}\log\tilde{g}-C_{10}\tilde{g},$
@@
for some $\phi(t)$ and $\psi(t)$ and constants $C_{8},C_{9},C_{10}$ to be determined. We deduce that the propagation of the new auxiliary function reads
$\displaystyle\,(\partial_{t}-\Delta_{f})F$
[…]
$\displaystyle$++ $\displaystyle\,\Big{(}\frac{M_{2}}{(1+t)^{2}}\phi+\psi^{\prime}+\frac{2M_{1}}{1+t}\psi+2\sigma\psi|\nabla\log\tilde{g}|^{2}+2\sigma C_{8}\log\tilde{g}+2\sigma C_{8}-\sigma C_{9}\Big{)}\frac{|\nabla\tilde{g}|^{2}}{\tilde{g}}.$
Recall the logarithmic gradient estimate
$\frac{\sigma t}{2M_{1}+1}|\nabla\log\tilde{g}|^{2}+\log\tilde{g}\leq C_{4}=\log 2M_{0}-\log\tau.$
Now we pick the coefficient function
$\psi(t)=\frac{t}{4M_{1}+2}$
to satisfy that
$2\sigma\psi|\nabla\log\tilde{g}|^{2}\leq\log 2M_{0}-\log\tau-\log\tilde{g}.$
We pick the other coefficient function
$\phi(t)=\frac{\sigma t^{2}}{(4M_{1}+M_{2}+2)(4M_{1}+2)}$
to cancel the first term on the right-hand side of the propagation of the auxiliary function $F$. Namely, we have
$\phi^{\prime}+\frac{4M_{1}+M_{2}}{1+t}\phi-\sigma\psi\leq\frac{2\sigma t+(4M_{1}+M_{2})\sigma t}{(4M_{1}+M_{2}+2)(4M_{1}+2)}-\frac{\sigma t}{4M_{1}+2}=0,$
and we deduce that
$\displaystyle(\partial_{t}-\Delta_{f})F\leq\Big{(}\frac{1}{2}+\frac{\sigma}{4M_{1}+2}+\log 2M_{0}-\log\tau+(2\sigma C_{8}-1)\log\tilde{g}-\sigma C_{9}\Big{)}\frac{|\nabla\tilde{g}|^{2}}{\tilde{g}}.$
The initial value of $F$ simply reads
$F(0,x)=-C_{8}\tilde{g}_{0}(\log\tilde{g}_{0})^{2}+C_{9}\tilde{g}_{0}\log\tilde{g}_{0}-C_{10}\tilde{g}_{0}.$
This allows us to choose $C_{8}=1/(2\sigma)$ to cancel the term involving $\log\tilde{g}$, and then choose
$C_{9}=\frac{1}{2\sigma}+\frac{1}{4M_{1}+2}+\frac{\log 2M_{0}}{\sigma}-\frac{\log\tau}{\sigma}=C_{9}^{\prime}-\frac{\log\tau}{\sigma}$
to deduce that $(\partial_{t}-\Delta_{f})F\leq 0$, and finally choose $C_{10}$ large enough to conclude that $F(0,\cdot)\leq 0$ as long as $\tau\geq 1$. The rest of the assumptions of Theorem 4.2 follow as in the short time case. Therefore, we arrive at $F\leq 0$, or
$\frac{\sigma t^{2}}{(4M_{1}+M_{2}+2)(4M_{1}+2)}\frac{|\nabla^{2}\tilde{g}|^{2}}{\tilde{g}^{2}}\leq\frac{(\log\tilde{g})^{2}}{2\sigma}-C_{9}\log\tilde{g}+C_{10}.$
Take $t=\tau/2$ and return to the original function $g$:
$\frac{|\nabla^{2}g|^{2}}{g}(\tau,m,x)\leq\frac{4(4M_{1}+M_{2}+2)(4M_{1}+2)}{\sigma\tau^{2}}\Big{(}\frac{(\log g)^{2}}{2\sigma}-(C_{9}^{\prime}-\log\tau)\log g+C_{10}\Big{)}.$ (4.19)
Combining (4.17) for $t\leq 1$ and (4.19) for $t\geq 1$, we can choose some appropriate large constants $B_{3},B_{4}$ to conclude the logarithmic Hessian estimate (4.2).
@@
Acknowledgments. X. F. and Z. W. are partially supported by the National Key R&D Program of China, Project Numbers 2021YFA1002800 and 2024YFA1015500, NSFC grant No.12171009, Young Elite Scientist Sponsorship Program by China Association for Science and Technology (CAST) No. YESS20200028 and the Fundamental Research Funds for the Central Universities (the start-up fund), Peking University.